Find e At where 2 5 A= -2 -4
Answer #1

Solution A= 2 5 1-2-4 let A= PD P1 where Dis Jordan matrix. AE ୧ Pept pt So first we find Jordan form of A. first we find eiglie 1,= -ti, 12 = -1-i Nars we find corresponding eigenvector for di = -lti (A-(-+)) -o =) [2+1-i 5 L-2 -4H-1 *49.7]C+7+[] T3P= (-3ti -3-1 00 ) 2 2 pe=1-112 1-31 % 1+3i 4 A= P -- py -Ito At = p/el-it P pt elt+i) = pet/e-it o (e-it eit Der leite cout